I have 9x-5=31 could you help me solve it step-by-step? Tomorrow I have my finals and I totally forgot how to do these. Thanks!
Answer by scott8148(5880)

add 5 to both sides to get 9x-5+5=31+5 ... 9x=36 ... divide by 9 to get 9x/9=36/9 ... x=4